OpenStudy (anonymous):

Medal and fan for help. Refer to the function g(x)=2+sqrt x+1. What is the domain of g(x)? @satellite73 @amistre64

OpenStudy (anonymous):

\[g \left( x \right)=2+\sqrt{x+1}\] \[\it is defined for all x+1\ge 0 or x \ge -1\] \[so domain is [-1,\infty )\]

OpenStudy (anonymous):

Thank you!!